Problem 11-1A Rogen Corporation manufactures a single product. The standard cost per unit of product is shown below Direct materials-1 pound plastic at $6 per pound Direct labor-0.50 hours at $12.35 per hour Variable manufacturing overhead Fixed manufacturing overhead 6.00 6.18 3.25 3.75 $19.18 Total standard cost per unit The predetermined manufacturing overhead rate is $14 per direct labor hour ($7.00 ÷ 0.50). It was computed from a master manufacturing overhead budget based on normal production of 2,500 direct labor hours (5,000 units) for the month. The master budget showed total variable costs of $16,250 ($6.50 per hour) and total fixed overhead costs of $18,750 $7.50 per hour). Actual costs for October in producing 4,400 units were as follows Direct materials (4,560 pounds) Direct labor (2,100 hours) Variable overhead Fixed overhead $27,816 26,355 21,558 11,142 $86,871 Total manufacturing costs The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignoredThe purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored. Compute all of the materials and labor variances. (Round answers to O decimal places, e.g. 125.) Total materials variance Materials price variance Materials quantity variance Total labor variance Labor price variance Labor quantity variance Compute the total overhead variance. Total overhead variance Click if you would like to Show Work for this question: Open Show Work LINK TO TEXT LINK TO TEXT VIDEO: SIMILAR PROBLEM
Answer #1

a.

Direct labor price variance = (SR – AR) x AH (12.35-12.55)*2100 = -420
Direct labor quantity variance = SR x (SH – AH) 12.35*(2500-2100) = 4940
Total direct labor variance = (SR x SH) – (AR x AH) (2500*12.35)-(2100*12.55) = 4520
=
Direct materials price variance = (SP – AP) x AQ (6-6.32)*4400 = -1408
Direct materials quantity variance = SP x (SQ – AQ) 6*(5000-4400) = 3600
Total direct materials variance = (SP x SQ) – (AP x AQ) (6*5000)-(6.32*4400) = 2192
where AR is Actual labour Rate (26355/2100)=12.55
SR is Standard Labour Rate
AH is actual hours
SH is Standard hours
SP is Standard Price
AP is Actual Price (27816/4400)= 6.32
SQ is Standard Quantity
AQ is Actual Quantity

b.

Calculation of Total Overhead Variance
1. Total Fixed overhead Variance = (Fixed overhead cost p.u*SQ-Actual Fixed Overhead) (3.75*5000)-11142      7,608 A
2. Total Variable overhead Variance (Variable overhead cost p.u*AQ-Actual Variable Overhead (3.25*4400)-21558     (7,258) B
Total Overhead Variance         350 A+B